# Artificial Intelligence in Image-based Cardiovascular Disease Analysis: A Comprehensive Survey and Future Outlook

### v3

@@ 4.2 Lifestyle and Environmental Factors
The analysis of population-based data illuminates the complexities and prevalence of cardiovascular conditions, highlighting the multifaceted influences on heart health (See. Figure 6 (left)). This includes genetic predispositions, lifestyle choices such as diet and physical activity, and environmental factors, all contributing to cardiovascular risk profiles [250, 251, 252]. Modern genomics connects genetic markers to heart health changes, while demographic factors like age and gender add another layer to cardiovascular risks. Integrating cardiovascular imaging with data on lifestyle, genetics, demographics, and environmental factors facilitates a holistic view of heart health, enabling tailored and more effective cardiovascular interventions [253].
Environmental influences, ranging from geographic location changes to lifestyle adaptations, air pollution, and social policies, significantly affect cardiovascular health [250, 251, 252]. The interplay between the natural environment, including aspects like sunlight exposure and green spaces, and the social environment, characterized by urbanization, pollution, and social networks, shapes lifestyle choices that impact cardiovascular risks. A comprehensive understanding of these interactions is crucial for devising innovative prevention and treatment strategies, addressing the global challenge of cardiovascular diseases. Advancements in AI and imaging genetics are poised to enhance our understanding of these complex relationships, offering promising avenues for research and clinical practice in cardiovascular health.
@@ 4.3 Interactions between Heart and Other Organs
The integration of AI in analyzing cardiovascular diseases through imaging techniques has significantly advanced our understanding of the heart’s connections with other bodily organs and systems. This comprehensive approach enables the identification of complex interrelations, such as the cardiovascular system’s impact on neurological health, and vice versa. For instance, AI algorithms are instrumental in analyzing cerebral blood flow and detecting stroke risks, providing insights into the cardiovascular origins of neurological events.
As shown in Figure 6 (right), a notable study [16] leveraging multiorgan MRI data from a large cohort has illuminated the phenotypic and genetic connections between cardiovascular health and brain attributes. This research has identified genomic loci that influence both heart and brain characteristics, suggesting a shared genetic foundation for cardiovascular and neurological conditions. Such findings underscore the potential of heart conditions to contribute causally to brain disorders, offering new perspectives on human health.
AI’s role extends beyond mere data analysis; it facilitates the extraction of critical insights from imaging data, encompassing the heart and its systemic interactions. By employing advanced AI techniques and multimodal data fusion, these models enhance the detection and quantification of abnormalities. This not only aids in accurate diagnosis and risk assessment but also enriches treatment strategies for cardiovascular diseases. The amalgamation of AI with extensive patient data, including imaging and clinical records, presents a holistic view of the cardiovascular system’s interplay with other organs, thereby promising to improve patient care and outcomes significantly.
5 Public Datasets
In recent years, several notable large-scale public image datasets for cardiovascular research have been made available, fostering collaboration and innovation in this domain.
Public datasets are invaluable for advancing AI in cardiac imaging by providing diverse, annotated data crucial for developing and testing algorithms. These resources facilitate collaborative efforts, benchmarking, and the sharing of innovative solutions across the research community, propelling the field forward. The accessibility of such datasets not only enhances model training and validation processes but also plays a pivotal role in the seamless integration of AI technologies into clinical settings, thereby elevating the standard of patient care.
The table below (Table I) categorizes available public datasets based on their focus areas and the dates they were made available. In the sections that follow, we delve into the specifics of selected datasets, offering insights into their composition, applications, and impact on cardiovascular disease research.
5.1 Dataset without Genetics Data
Automatic Cardiac Diagnosis Challenge (ACDC) [233]. The ACDC dataset is tailored for benchmarking cardiac diagnosis and segmentation algorithms, featuring cardiac MRI scans from diverse cardiovascular conditions, annotated for left ventricle contours, thus facilitating the development of precise AI segmentation and diagnostic tools.
Multi-Ethnic Study of Atherosclerosis (MESA) [254]. The Multi-Ethnic Study of Atherosclerosis (MESA) dataset, renowned in cardiovascular research, includes MRI scans from a broad demographic, paired with extensive clinical data, enabling in-depth studies on conditions like coronary artery disease and aiding in the advancement of AI diagnostic algorithms.
Cardiac Acquisitions for Multi-structure Ultrasound Segmentation (CAMUS) [24]. The CAMUS contains 2D echocardiographic sequences with two and four-chamber views of 500 patients. The size of this dataset and its tight connection to every-day clinical issues give the possibility to train AI methods to automatically analyze echocardiographic data. In addition, CAMUS includes manual expert annotations for the left ventricle endocardium (LVEndo), the myocardium (epicardium contour more specifically, named LVEpi) and the left atrium (LA).
Aortic Stiffness 4D Flow (AS4DF) [245]. The AS4DF data includes MRI data of a subject-specific thoracic aorta model embedded into a physiological flow circuit. The aorta model was manufactured using compliant 3D-printing technology. Multiple imaging experiments were performed on multiple models that differed regarding wall compliance. The acquired data includes 2D phase contrast, 2D gradient echo cine, 3D spoiled gradient echo, and 4D flow (at three temporal resolutions).
Coronary Artery Calcium (CAC) [255]. The CAC datasets consist of CT scans of the coronary arteries, focusing on the identification and quantification of coronary artery calcification. These datasets provide valuable information for assessing cardiovascular risk and diagnosing coronary artery disease. They facilitate the development of AI algorithms for automated calcium scoring and the prediction of cardiovascular events.
Cardiac Atlas Project (CAP) [256]. CAP is a global consortium and an online platform dedicated to the integration and sharing of cardiac imaging examinations. It encompasses parametric model-based functional analyses and related clinical data. The initiative has led to the creation of an extensive database featuring a wide range of cardiac imaging examinations, associated models, and derived measurements. The primary aim is to develop a comprehensive, web-accessible atlas that delineates both the normal and diseased states of the heart’s structure and function, serving clinical, research, and educational needs [241].
CardioVascular Research Grid (CVRG) [242]. CVRG is a collaborative platform designed to advance cardiovascular research by facilitating the sharing and integration of data across the field. It provides researchers and clinicians with access to an extensive database, encompassing a wide range of cardiovascular data including imaging studies, clinical information, and research findings. This platform not only accelerates the understanding and treatment of heart diseases but also contributes to the development of personalized medicine in cardiology. By centralizing resources and fostering a global research community, the CVRG plays a crucial role in informing clinical practices and advancing global health initiatives related to cardiovascular diseases [257].
Stanford Center for Artificial Intelligence in Medicine and Imaging (AIMI) shared dataset [258]. The AIMI dataset offers a repository of clinical imaging data, including echocardiograms and cardiac MRI data, curated for research purposes. These datasets support the development of AI tools for cardiovascular research.
5.2 Dataset with Genetics Data
UK Biobank (UKB) [259]. The UKB stands as a monumental biomedical database, offering genetic and health data from half a million UK participants, making it a cornerstone for research on prevalent and life-threatening diseases. It includes cardiac MRI among other imaging data, making it a treasure trove for AI-driven cardiovascular research.
Heart Imaging Genetics Knowledge Portal (Heart-KP) [260]. Heart-KP is a powerful resource for advancing cardiovascular research by integrating heart imaging and genetic data. This portal provides GWAS results for heart imaging traits, including data on left and right ventricle traits, atrium traits, and aorta traits. It supports extensive genetic and phenotypic data analysis for heart-related research. It enables researchers to explore the genetic basis of heart structure and function, contributing to the development of new diagnostic and therapeutic strategies for heart diseases.
American Heart Association (AHA) Data Repositories [258]. The AHA repositories include datasets that provide detailed imaging data for cardiovascular research. These can include MRI, CT scans, echocardiograms, and other types of heart imaging studies. The repositories also offer genetic data, including genome-wide association studies (GWAS), exome sequencing, and other genetic information relevant to cardiovascular health. This data helps in understanding the genetic underpinnings of heart diseases and conditions, which archives and distributes results of studies investigating the interaction of genotype and phenotype, including those related to heart disease.
Figure 7: Exploring the future with AI Foundation Models: Harnessing medical imaging as the core modality alongside auxiliary data for clinical innovation and combating cardiovascular diseases through precision medicine and digital twin technologies.
@@ 6 Future Directions
While significant progress has been made in leveraging AI for image-based CVD analysis, there are still several avenues for future research and development. We discuss several promising directions, and the overview is illustrated in Figure 7.
Precision Medicine in Cardiology. Precision medicine is revolutionizing healthcare by customizing medical treatments to each individual’s unique genetic makeup, environmental factors, and lifestyle. This approach has the potential to dramatically transform how we diagnose, track, and treat diseases [261]. A particularly promising area is the application of AI in integrating heart medical imaging with other modalities, like genomics, proteomics, and metabolomics, to improve clinical decision-making and disease management [262]. This integration offers a more comprehensive understanding of a patient’s health status, but also raises challenges such as data privacy, ethical considerations in the use of personal health information, and the need for stringent validation of AI predictions to ensure reliability and safety [263].
@@
Digital Twins (DT) of Heart. The goal of creating a digital twin of the heart is to facilitate tailored and precise treatments for cardiovascular diseases [264]. The process involves collecting a wide array of patient data, including advanced imaging, genetic profiles, and physiological metrics [265]. These data points are synthesized using sophisticated AI techniques to construct a dynamic, virtual model of the patient’s heart, reflecting its detailed anatomy and functions [266]. The DT model enables simulations to test various therapeutic strategies, offering insights into the most effective treatments. It’s continuously updated with new patient data, allowing for real-time health monitoring and enabling informed clinical decisions to adjust treatments as needed. Digital twins of the heart offer significant benefits by enabling healthcare professionals to test and refine treatment strategies in a risk-free environment [267]. This innovative approach can lead to safer patient care, lower healthcare costs, and improved outcomes. Additionally, digital twins have the potential to spur new therapeutic discoveries and enhance our understanding of cardiovascular diseases [268]. However, their widespread implementation faces hurdles such as ensuring data privacy, meeting computational demands, and achieving model accuracy.
@@
AI-based Cardiac Image Generation. Digitally Reconstructed Radiographs (DRR) serve as a pivotal technique in medical imaging, simulating X-ray images from 3D medical data such as CT or MRI scans, thereby reducing unnecessary radiation exposure for patients [269]. The integration of advanced deep learning models, including 3D convolutional neural networks, GANs [70], variational autoencoders [270], and diffusion model [271, 119] enhances the reconstruction quality but demands considerable computational resources and extensive training data [272]. The advent of Neural Radiance Fields (NeRF) has brought significant advancements in natural image reconstruction, though their application in medical imaging remains challenging due to the complexity of medical data [273]. Innovations like Medical image NeRFs [274, 275] are making strides by learning to map radiance values to pixels, revealing intricate details of internal anatomy from 2D images, and offering potential for reducing radiation risks and examination costs, especially in orthopedics and surgery.
Explainable CVDs Diagnosis Models with LLMs. Large Language Models (LLMs) hold significant promise in revolutionizing the diagnosis of CVDs, the leading global cause of mortality. By leveraging their vast data processing capabilities, LLMs can substantially improve diagnostic precision. However, the opaque nature of these models raises concerns about transparency and trust among medical professionals and patients [276]. To address this, it’s crucial to incorporate explainability into LLMs, ensuring they not only deliver accurate diagnoses but also provide comprehensible rationales for their decisions [277, 278, 279]. Enhancing the explainability of LLMs [280] involves creating or improving interpretability methods tailored to medical diagnostic needs. This step is essential for the models’ acceptance and practical application in healthcare. Moreover, thorough real-world clinical validations are imperative to confirm the model’s dependability and effectiveness across various healthcare applications.
AI Foundation Models for Heart Health. AI is increasingly becoming the primary approach for analyzing CVD through imaging [281], aiming to reduce costs and the reliance on invasive procedures [203, 199, 196]. AI algorithms have been developed for tasks such as disease classification, risk prediction, treatment planning, and clinical decision-making in cardiovascular imaging [282]. However, most current AI approaches in this field rely on task-specific models [79], which may not fully capture the complexity and diversity of medical data due to their smaller size and task-focused design.
@@
The emergence of AI foundation models, particularly prominent in computer vision and natural language processing [283, 284, 285, 286, 287], represents a significant shift. These large-scale models, with millions or billions of parameters, are capable of handling more complex data and identifying a wide range of patterns and relationships, often achieving state-of-the-art results across various tasks due to their scalability and generalization capabilities [288, 289, 290, 291, 292]. In cardiovascular health, foundation models have the potential to be adapted or fine-tuned for specific CVD analysis tasks [293], such as Atrial Fibrillation [20], Ischemic Heart Disease [91], and Coronary Artery Diseases [203]. Adapting these models to the diverse imaging modalities used in CVD analysis presents challenges due to the significant differences from natural images, based on various physics-based properties and energy sources [294]. A potential solution involves training foundation models on both medical and natural images through fine-tuning approaches [295, 296], aiming to provide a robust foundational solution for clinical challenges in CVD. This advancement could significantly improve the effectiveness and efficiency of diagnosing and treating cardiovascular diseases, marking a substantial step forward in the field [297]. Furthermore, developing more efficient neural network architectures is crucial for advancing foundation models. Notably, Mamba [298], a State Space Model, has emerged as a significant innovation for its adeptness at efficiently modeling long-range dependencies in sequential data, attributed to its exceptional memory efficiency and computational speed [299, 300].
@@ 7 Conclusion
The integration of AI in cardiovascular imaging is significantly enhancing diagnostic accuracy, personalized treatment planning, and early disease detection in the field of cardiology. This paper offers an in-depth examination of AI applications in cardiovascular disease analysis through imaging, covering a range of topics from imaging modalities and processing techniques to disease-specific analysis methods. Challenges such as the collection, annotation, and standardization of large-scale datasets are discussed, emphasizing the critical role of public datasets and code repositories in promoting research collaboration and algorithm validation. The paper also sheds light on promising future directions, including imaging genetics, the concept of digital twins, and the utilization of AI foundational models, which are expected to drive substantial progress in cardiovascular imaging. Looking ahead, the paper anticipates further advancements in AI algorithms, hardware, and imaging technologies, which will lead to more precise and efficient diagnostic and treatment processes. Moreover, the integration of AI with electronic health records, genomics, and other clinical data will facilitate a comprehensive and personalized approach to cardiovascular care, aligning with the evolving landscape of patient-centered healthcare.
Acknowledgment
Copyrights of the figures belong to their own authors and/or holders (sources are cited in the captions). The author sincerely thanks all readers for taking the time to read the manuscript. If any related work has been overlooked, please do not hesitate to contact us. The feedback will be greatly appreciated and help us considerably improve this survey.
Xin Wang is supported by SUNY at Albany Start-up Grant. The research of Dr. Hongtu Zhu is partially supported by the National Institute on Aging (NIA) of the National Institute of Health (NIH) under Award Number RF1AG082938, U01AG079847, and R01AR082684.

### v4

@@ 4.2 Lifestyle and Environmental Factors
The analysis of population-based data illuminates the complexities and prevalence of cardiovascular conditions, highlighting the multifaceted influences on heart health. This includes genetic predispositions, lifestyle choices such as diet and physical activity, and environmental factors, all contributing to cardiovascular risk profiles [munzel2022environmental, bhatnagar2017environmental, powell2022social]. Modern genomics connects genetic markers to heart health changes, while demographic factors like age and gender add another layer to cardiovascular risks. Integrating cardiovascular imaging with data on lifestyle, genetics, demographics, and environmental factors facilitates a holistic view of heart health, enabling tailored and more effective cardiovascular interventions [barratt2023environmental]. Environmental influences, ranging from geographic location changes to lifestyle adaptations, air pollution, and social policies, significantly affect cardiovascular health [munzel2022environmental, bhatnagar2017environmental, powell2022social]. The interplay between the natural environment, including aspects like sunlight exposure and green spaces, and the social environment, characterized by urbanization, pollution, and social networks, shapes lifestyle choices that impact cardiovascular risks. A comprehensive understanding of these interactions is crucial for devising innovative prevention and treatment strategies, addressing the global challenge of cardiovascular diseases. Advancements in AI and imaging genetics are poised to enhance our understanding of these complex relationships, offering promising avenues for research and clinical practice in cardiovascular health.
@@ 4.3 Interactions between Heart and other Organs
The integration of AI in analyzing cardiovascular diseases through imaging techniques has significantly advanced our understanding of the heart’s connections with other bodily organs and systems. This comprehensive approach enables the identification of complex interrelations, such as the cardiovascular system’s impact on neurological health, and vice versa. For instance, AI algorithms are instrumental in analyzing cerebral blood flow and detecting stroke risks, providing insights into the cardiovascular origins of neurological events. A notable study [zhao2023heart] leveraging multiorgan MRI data from a large cohort has illuminated the phenotypic and genetic connections between cardiovascular health and brain attributes. This research has identified genomic loci that influence both heart and brain characteristics, suggesting a shared genetic foundation for cardiovascular and neurological conditions. Such findings underscore the potential of heart conditions to contribute causally to brain disorders, offering new perspectives on human health. AI’s role extends beyond mere data analysis; it facilitates the extraction of critical insights from imaging data, encompassing the heart and its systemic interactions. By employing advanced AI techniques and multimodal data fusion, these models enhance the detection and quantification of abnormalities. This not only aids in accurate diagnosis and risk assessment but also enriches treatment strategies for cardiovascular diseases.
4.4 Section Summary
In summary for section 4, we discussed how large-scale imaging data is leveraged to understand CVD trends, risk factors, and progression patterns across diverse populations. Population-level imaging studies can identify associations between anatomical variations and CVD prevalence, uncovering insights into how risk factors (such as age, gender, ethnicity, and lifestyle) impact disease development. By analyzing imaging data from broad, heterogeneous cohorts, these studies provide normative baselines and detect atypical presentations, which are invaluable for risk stratification and early intervention. Additionally, population-based CVD analysis informs public health policies, enabling targeted prevention efforts and resource allocation tailored to the cardiovascular health needs of specific communities or demographics.
@@ 5 Future Directions
Exploring the future of cardiovascular health through AI-driven models anchored in medical imaging, alongside auxiliary data such as genetics, clinical history, and lifestyle factors, opens a path to transformative clinical innovation. Medical imaging provides a foundational view of the heart’s structure and function, and when combined with auxiliary data, AI technologies can deliver unprecedented insights into cardiovascular disease (CVD) that extend beyond traditional analysis. This integrative approach enables personalized, precise diagnosis and treatment strategies tailored to each patient’s unique profile, making precision medicine a reality in cardiovascular care.
5.1 Digital Twins (DT) of Heart
@@
The goal of creating a digital twin of the heart is to facilitate tailored and precise treatments for CVD [corral2020digital]. The process involves collecting a wide array of patient data, including advanced imaging, genetic profiles, and physiological metrics [tao2018digital]. These data points are synthesized using sophisticated AI techniques to construct a dynamic, virtual model of the patient’s heart, reflecting its detailed anatomy and functions. The DT model enables simulations to test various therapeutic strategies, offering insights into the most effective treatments. Digital twins of the heart offer significant benefits by enabling healthcare professionals to test and refine treatment strategies in a risk-free environment [coorey2022health]. This innovative approach can lead to safer patient care, lower healthcare costs, and improved outcomes. Additionally, digital twins have the potential to spur new therapeutic discoveries and enhance our understanding of cardiovascular diseases [ghatti2023digital]. However, their widespread implementation faces hurdles such as ensuring data privacy, meeting computational demands, and achieving model accuracy.
5.2 AI-based Cardiac Image Generation
@@
Digitally Reconstructed Radiographs (DRR) serve as a pivotal technique in medical imaging, simulating X-ray images from 3D medical data such as CT or MRI scans, thereby reducing unnecessary radiation exposure for patients [lo2012extraction]. The integration of advanced deep learning models, including 3D convolutional neural networks, GANs [goodfellow2020generative], variational autoencoders [kingma2019introduction], and diffusion model [he2023dmcvr, kazerouni2023diffusion] enhances the reconstruction quality but demands considerable computational resources and extensive training data [maken20232d]. The advent of Neural Radiance Fields (NeRF) has brought significant advancements in natural image reconstruction, though their application in medical imaging remains challenging due to the complexity of medical data [mildenhall2021nerf]. Innovations like Medical image NeRFs [corona2022mednerf, hu2023umednerf] are making strides by learning to map radiance values to pixels, revealing intricate details of internal anatomy from 2D images, and offering potential for reducing radiation risks and examination costs, especially in orthopedics and surgery.
5.3 AI Foundation Models for Heart Health
@@
AI is increasingly becoming the primary approach for analyzing CVD through imaging [olawade2024advancements], aiming to reduce costs and the reliance on invasive procedures [gao2020learning, wu2019automated, kong2020learning]. AI algorithms have been developed for tasks such as disease classification, risk prediction, treatment planning, and clinical decision-making in cardiovascular imaging [ma2019iterative]. However, most current AI approaches in this field rely on task-specific models [lu2023attention], which may not fully capture the complexity and diversity of medical data due to their smaller size and task-focused design. The emergence of AI foundation models, particularly prominent in computer vision and natural language processing [zhou2023comprehensive, sun2023survey, zhou2023foundation, li2023multimodal], represents a significant shift. These large-scale models, with millions or billions of parameters, are capable of handling more complex data and identifying a wide range of patterns and relationships, often achieving state-of-the-art results across various tasks due to their scalability and generalization capabilities [radford2018improving, nguyen2023lvm, wang2023medfmc, wang2023foundation, lin2024robust]. In cardiovascular health, foundation models have the potential to be adapted or fine-tuned for specific CVD analysis tasks, such as Atrial Fibrillation [li2022medical], Ischemic Heart Disease [xue2020automated], and Coronary Artery Diseases [gao2020learning]. Adapting these models to the diverse imaging modalities used in CVD analysis presents challenges due to the significant differences from natural images, based on various physics-based properties and energy sources [zhang2023challenges]. A potential solution involves training foundation models on both medical and natural images through fine-tuning approaches [li2023llava, ma2024segment], aiming to provide a robust foundational solution for clinical challenges in CVD. This advancement could significantly improve the effectiveness and efficiency of diagnosing and treating CVD, marking a substantial step forward in the field [wang2024u].
5.4 Explainable CVDs Diagnosis Models with LLMs
Large Language Models (LLMs) hold significant promise in revolutionizing the diagnosis of CVDs, the leading global cause of mortality. By leveraging their vast data processing capabilities, LLMs can substantially improve diagnostic precision [liu2025medchat]. However, the opaque nature of these models raises concerns about transparency and trust among medical professionals and patients. To address this, it’s crucial to incorporate explainability into LLMs, ensuring they not only deliver accurate diagnoses but also provide comprehensible rationales for their decisions [teng2022survey]. Enhancing the explainability of LLMs [zhao2023explainability] involves creating or improving interpretability methods tailored to medical diagnostic needs. This step is essential for the models’ acceptance and practical application in healthcare. Moreover, thorough real-world clinical validations are imperative to confirm the model’s dependability and effectiveness across various healthcare applications.
@@ 6 Conclusion
The integration of AI in cardiovascular imaging is significantly enhancing diagnostic accuracy, personalized treatment planning, and early disease detection in the field of cardiology. This paper offers an in-depth examination of AI applications in cardiovascular disease analysis through imaging, covering a range of topics from imaging modalities and processing techniques to disease-specific analysis methods. Challenges such as the collection, annotation, and standardization of large-scale datasets are discussed, emphasizing the critical role of public datasets and code repositories in promoting research collaboration and algorithm validation. The paper sheds light on promising future directions, including imaging genetics, digital twins, and the utilization of AI foundational models, which are expected to drive substantial progress in cardiovascular imaging. Moreover, the integration of AI with electronic health records, genomics, and other clinical data will facilitate a comprehensive and personalized approach to cardiovascular care, aligning with the evolving landscape of patient-centered healthcare.
7 ACKNOWLEDGMENTS
Copyrights of the figures belong to their own authors and/or holders (sources are cited in the captions). Dr. Zhu’s work was partially supported by the Gillings Innovation Laboratory on generative AI and by grants from the National Institute on Aging (NIA) of the National Institutes of Health (NIH), including 1R01AG085581, and RF1AG082938, the National Institute of Mental Health (NIMH) grant 1R01MH136055, and the NIH grants R01AR082684, and 1OT2OD038045-01. The content of this paper is solely the responsibility of the authors and does not necessarily represent the official views of these institutions.
8 DISCLOSURE STATEMENT
The authors are not aware of any affiliations, memberships, funding, or financial holdings that might be perceived as affecting the objectivity of this review.